( Payback period Calculations) You are considering three independent projects, project A, project B, and project C. Given the following cash flow information, calculate the payback period for each. if you require a 3-year payback before an investment can be accepted, which project(s) would be accepted?

Project A Project B Project C
Initial Outlay -$1,100 -$9,000 -$6,000
Inflow year 1 600 4,000 2,000
inflow year 2 300 3,000 2,000
inflow year 3 200 3,000 3,000
inflow year 4 100 3,000 3,000
inflow year 5 500 3,000 3,000

1) What is the payback period of project A?

2) What is the payback period of project B?

3) What is the payback period of project C?

Which project(s) would be accepted
Answer #1

1) Payback period formula Project A = Years before recovery + Cost not covered in that year/ Cash flow for that year =3
Because in year 3 Sum of cash flows= 600+300+200 =11000

2) Payback period formula Project B = Years before recovery + Cost not covered in that year/ Cash flow for that year
=2+(9000-7000)/3000 =2.67

3) Payback period formula Project C = Years before recovery + Cost not covered in that year/ Cash flow for that year
=2+(6000-2000-2000)/3000 =2.67

Project B, Project C should be accepted because Payback Period is less than 3.